User Tools

Site Tools


handover:rd1502

RD1502

Hobart 26m

Module: HOB+0107/16000 (Backup: BKG+0142/4000)

Data volume at beginning: 8318

  1. Fs complained 've' was not in remote at start of experiment. Switching to remote resolved.
  2. Attenuation recommended by notes was too low. Changed to 6db in IFD as per CRDS73 prc file.
  3. 02:15 UT: He supply pressure has shot up from <285 psi to 310 psi because Brett Reid introduced new supply (Ross).
You could leave a comment if you were logged in.
/home/www/auscope/opswiki/data/pages/handover/rd1502.txt · Last modified: 2015/04/10 12:57 by liza